40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Jul 23 22:34:07 2008, Yoichi, Update, LSC, Arm cavity g-factor measurement 
    Reply  Thu Jul 24 01:27:00 2008, Koji, Update, LSC, Arm cavity g-factor measurement 
       Reply  Thu Jul 24 02:57:26 2008, rob, Update, LSC, Arm cavity g-factor measurement 
          Reply  Thu Jul 24 08:09:26 2008, Yoichi, Update, LSC, Arm cavity g-factor measurement 
Message ID: 728     Entry time: Wed Jul 23 22:34:07 2008     Reply to this: 730
Author: Yoichi 
Type: Update 
Category: LSC 
Subject: Arm cavity g-factor measurement 
I tried the same thing as the X-arm to the Y-arm.
I'm puzzled. I found exactly the same behavior as the X-arm in the AS166 demodulated signals, whereas I expected different resonance frequency because of the arm length difference.

Here is more detailed account of the measurement today.

I locked the Y-arm and mis-aligned the end mirror in Yaw until the transmission power gets half.
Then I injected a 30Hz sinusoid into the error point of the Y-arm servo to shake the ETMY.
I observed AS166_I and AS166_Q as I changed the 166MHz frequency.

At 165.977MHz, both AS166_I and AS166_Q showed the 30Hz signal (15cnt p-p).
At 165.981MHz, Only I phase showed the 30Hz signal (40cnt p-p). No signal in Q.
At 165.984MHz, I and Q became the same amplitude again (20cnt p-p).
At 165.987MHz, Only Q phase showed the 30Hz signal (40cnt p-p). No signal in I.

Outside the above range, the signal decreases as the frequency go away. I think this is (at least partly) because the 166MHz sidebands no longer go through the MC at those frequencies.

I then locked the X-arm to the TEM01 mode. I saw exactly the same behavior as described above. This could be the resonance of TEM02 mode. I was expecting to see the resonance of TEM00 mode at the opposite side, but nothing there.

I unlocked the arm cavities and tried the same frequency scan of the 166MHz with one of the end mirrors shaken at 30Hz. I saw no signal at the AS166 port.
I also tried locking Y-arm and shaking the ETMX. No signal.
So it has to be something to do with the cavity resonance.

Since the MC transmission curve for 166MHz is folded in the measurement, it makes the interpretation of the results harder.
ELOG V3.1.3-